Using FOIL, simplify the expression "(3x + 2)(3x - 2)." Show that a particular factoring formula leads to the same answer. Thanks!
Answer by rapaljer(3625)

(3x+2)(3x-2)

F=3x*3x=9x^2
0=3x*-2 =-6x
I=2*3x=6x
L=+2*-2=-4

F+O+I+L=9x^2 -6x+6x-4
F+O+I+L= 9x^2 -4

This actually ends up the DIFFERENCE of TWO SQUARES.

Dr. Robert J. Rapalje, Retired